OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2 I think I know the answer but at the same time I feel i'm wrong... Anyone willing to help?

OpenStudy (anonymous):

Okay well the formula is Y=mx+b where m = slope and b = y intercept so whats do you think ?(:

OpenStudy (anonymous):

I think it's y=1/3x-2 because you plug 1/3 into for m and -2 for b I didn't know where the y intercept went for sure but now I think I know :3

OpenStudy (anonymous):

if the slope is 3 why do you think its 1/3 ?

OpenStudy (anonymous):

Because if it's just 3 isn't it automatically 1/3 as the slopee?

OpenStudy (texaschic101):

hes thinking of perpendicular lines ?

OpenStudy (anonymous):

No i dont believe so unless you want the inverse for perpendicular lines

OpenStudy (anonymous):

Take a look at this - http://www.regentsprep.org/Regents/math/geometry/GCG1/EqLines.htm

OpenStudy (anonymous):

then it would be negative :/

OpenStudy (anonymous):

Wait so when it's not a fraction for exmaple just the number 4 then it would be 4/1?

OpenStudy (texaschic101):

true...forgot the sign

OpenStudy (anonymous):

Yes i believe so

OpenStudy (anonymous):

yep 4/1 = 4

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Oh man I mixed that up. Oh well I understand now everyone. ;D Thanks a bunch

OpenStudy (anonymous):

No problem !

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (anonymous):

gracis lol